# Fully Spiking Neural Network for Legged Robots

### v1

@@ IV-B2 Cassie
The training results of Cassie robot have proven to be highly successful, as evidenced by the findings depicted in Figure 6. The stability observed in the angular velocity following along the yaw axis is particularly noteworthy. This stability is crucial in ensuring effective control and balance while navigating diverse terrains. Furthermore, the robot’s impressive ability to reach the highest level of terrain height, as indicated by a reward score of up to $6$ (with each unit representing a layer in the pyramid ladder), showcases its adaptability in conquering rugged landscapes. These findings highlight the Cassie robot’s capacity to successfully traverse challenging terrains under the high control frequency of SNNs while maintaining stability and elegance. Consequently, the feasibility of our approach is substantiated, affirming its potential deployment in complex and ever-changing real-world scenarios.
@@ IV-B3 MIT Humanoid
In the videos presented, the robots trained using SNN-based policies exhibit remarkable characteristics, including smooth, stable, and natural gaits. Whether it is the agile traversal of challenging terrains by A1 and Cassie or the unrestricted running of the MIT Humanoid, the performance showcased by our approach is undeniably superior. Notably, the low energy consumption associated with SNN further positions it as a promising alternative to ANN. With its ability to achieve exceptional performance and energy efficiency, SNN holds great potential in the field of robotics.
V Conclusion and Limitation
This study presents the integration of PopSAN with two cutting-edge reinforcement learning techniques, namely history trajectory and imitation learning. This innovative approach enables SNNs to achieve performance comparable to ANNs. The successful training of SNN-based policy networks using these methods highlights the versatility of SNNs in policy gradient-based DRL algorithms, thereby opening up new horizons for their application in various reinforcement learning tasks, including continuous control.
Additionally, the study explores the suitability of SNNs for high-frequency control tasks, demonstrating that agents powered by SNNs can devise exceptional strategies even at an elevated inference frequency of 500Hz. These findings emphasize the advantages of SNN-based control methods over ANNs, particularly in scenarios with limited computational resources. They showcase the remarkable energy efficiency and robustness of SNNs, highlighting their potential in various applications.
However, due to the difficulty in obtaining neuromorphic chips, our experiments could not be conducted on a physical machine, consequently slightly diminishing the strength of our argument regarding the energy efficiency of SNNs.
@@ V Conclusion and Limitation
Acknowledgment
The authors would like to express their gratitude to Jingtong Ma and Chao Chen for their valuable contributions and insightful discussions.

### v2

@@ IV-B2 Cassie
In Cassie’s experiments, a trapezoidal pyramid environment was utilized, with terrain level indicates the absolute value of the robot’s elevation displacement (a value of 6 signifies reaching the top of the sixth-order pyramid). The training results of the Cassie robot have proven to be highly successful, as shown in Figure 6. The stability of the angular velocity along the yaw axis is notable. This stability is crucial for maintaining effective control and balance while navigating diverse terrains. Furthermore, the robot’s impressive ability to reach the highest terrain levels demonstrates its adaptability in conquering rugged landscapes. These findings highlight the Cassie robot’s ability to navigate difficult terrains effectively using SNNs at a high control frequency while maintaining stability. Consequently, the feasibility of our approach is confirmed, supporting its potential deployment in complex and dynamic real-world scenarios.
@@ IV-B3 MIT Humanoid
The performance demonstrated by our approach, whether through A1 and Cassie’s agile traversal of challenging terrains or the MIT Humanoid’s unrestricted running, is undeniably superior.
IV-C Continuity Comparison
Compared to ANN, SNN also shows stronger robustness. This is because in SNNs the signaling relies on thresholding, leading to the direct exclusion of some noise. The stochastic behavior of spiking neuron dynamics also enhances the robustness of the network to external noise. We added Gaussian noise to the robot’s movement commands, and tested the performance of the networks with two strategies, ANN and SNN, for sigma of 0.1, 0.2, and 0.3, respectively, using the following of the robot’s y-axis linear velocity as an index. The experiment (Figure 8) proves that as the noise increases, the SNN’s ability to resist noise becomes more and more obvious, and its measurements become closer to the desired value and are accompanied by less fluctuations. On the other hand, the situation of ANN is not optimistic, at sigma of 0.3, the trained strategy is completely insufficient to support the robot to walk normally, and falls occurred. Under the same noise interference, the robot with SNN as the policy network can still walk smoothly.
Figure 8: Linear velocity following case. The original command was set to 0.4, and Gaussian noise with a scaling factor of 0.1 was introduced. SNN consistently surpasses that of ANN when subjected to equivalent levels of noise, with this difference widening as the noise level intensifies. At a sigma of 0.3 for Gaussian noise, the robot with SNN as its strategy maintained stable command following and smooth walking, whereas the ANN version of the robot experienced instability and tumbled.
IV-D Estimation of Energy Consumption
As mentioned earlier, one of the primary advantages of our SNN policy is its minimal energy usage. The assessment of energy consumption in a SNN is complex because the floating-point operations (FLOPs) in the initial encoder layer are MAC, whereas all other Conv or FC layers are AC. Building upon prior research[43, 44, 45, 46] conducted by SNN, it is assumed that the data utilized for various operations is represented in 32-bit floating-point format in 45nm technology[45], with $E_{MAC}$ = 4.6pJ and $E_{AC}$ = 0.9pJ. The energy consumption equations simulated for SNN are provided as follows:
$E_{model}=$ $E_{MAC}\cdot FL^{1}_{SNNConv}+$ (4)
$E_{AC}\cdot(\sum_{n=2}^{N}FL^{n}_{SNNConv}+\sum_{m=1}^{M}FL^{m}_{SNNFC})$
Experimental results (Table II) demonstrate that our approach offers a significant energy efficiency improvement compared to the conventional ANN architecture. Specifically, it achieves energy savings of $95.65\%$, $81.49\%$, and $60.34\%$ at T = 1, 2 and 3, respectively.
TABLE II: Energy Comparison(×$\times 10^{-6}$ mJ)
Method Actor(T=1) Actor(T=2) Actor(T=3)
ANN Model 85.96 85.96 85.96
SNN Model(ours) $3.35$ $15.36$ $35.22$
Energy Saving 96.10% 82.13% 59.03%
V Conclusion
This study presents the integration of SNN with history trajectory and imitation learning, which achieves performance comparable to ANNs, highlights the versatility of SNNs in policy gradient-based DRL algorithms. This opens up new horizons for application of SNNs in various reinforcement learning tasks, including continuous control.
Additionally, our approach offers significant advantages in energy efficiency, addressing signal noise, and high-frequency control. These advantages are significant for improving structural integrity and robustness in practical situations. It also has potential to reduce costs related to robot development and enables the implementation of advanced sensing systems in robotic platforms.